# APRIL: Active Partial Rollouts in Reinforcement Learning to tame long-tail generation

### v1

@@ Abstract
Reinforcement learning (RL) has become a cornerstone in advancing large-scale pre-trained language models (LLMs). Successive generations—including GPT-o series, DeepSeek-R1, Kimi-K1.5, Grok 4, and GLM-4.5—have relied on large-scale RL training to enhance reasoning and coding capabilities. To meet the community’s growing RL needs, numerous RL frameworks have been proposed. Most of these frameworks primarily rely on inference engines for rollout generation and training engines for policy updates. However, RL training remains computationally expensive, with rollout generation accounting for more than 90% of total runtime. Besides, its efficiency is often constrained by the long-tail distribution of rollout response lengths, where a few lengthy responses stall entire batches, leaving GPUs idle and underutilized. As model and rollout sizes continue to grow, this bottleneck increasingly limits scalability. To address this challenge, we propose Active Partial Rollouts in Reinforcement Learning (APRIL), which mitigates long-tail inefficiency. In the rollout phase, APRIL over-provisions rollout requests, terminates once the target number of responses is reached, and recycles incomplete responses for continuation in future steps. This strategy ensures that no rollouts are discarded while substantially reducing GPU idle time. Experiments show that APRIL improves rollout throughput by at most 44% across commonly used RL algorithms (GRPO, DAPO, GSPO), accelerates convergence, and achieves at most 8% higher final accuracy across tasks. Moreover, APRIL is both framework- and hardware-agnostic, already integrated into the slime111slime repository: https://github.com/THUDM/slime RL framework, and deployable on NVIDIA and AMD GPUs alike. Taken together, this work unifies system-level and algorithmic considerations in proposing APRIL, with the aim of advancing RL training efficiency and inspiring further optimizations in RL systems. Our codebase is available at: https://github.com/RLsys-Foundation/APRIL
@@ 1 Introduction
This need has been actively addressed by the open-source community, resulting in scalable RL training frameworks such as OpenRLHF (Hu et al., 2025), verl (Sheng et al., 2025), Areal (Fu et al., 2025), and slime (THUDM, 2025). These frameworks typically integrate advanced distributed engines to enhance efficiency—employing vLLM (Kwon et al., 2023) or SGLang (Zheng et al., 2024) as inference backends for LLM rollouts, FSDP (Zhao et al., 2023) or Megatron-LM (Shoeybi et al., 2019) as training backends for LLM optimization, and Ray (Moritz et al., 2018) for orchestrating parallel training and inference to maximize resource utilization. Despite extensive infrastructure-level optimizations, the rollout phase remains the dominant bottleneck: each input requires the LLM to generate $\mathcal{N}$ responses of length up to $\mathcal{L}$ in an auto-regressive manner, which typically accounts for over 90% of the total RL training runtime (Hu et al., 2025). Besides, the response length of the rollouts for each input instance often varies significantly in most tasks (3.2). As a result, in any given training batch, a small number of examples may require exceptionally long responses. This ”long-tail” distribution causes the entire batch to stall until the longest-running instance rollout is completed, leading to severe inefficiency and under-utilization of computational resources, as faster-generating workers remain idle (Zhong et al., 2025c). Moreover, with recent findings and the emerging trend of scaling up rollout inference ($\mathcal{N}$ and $\mathcal{L}$) to enhance LLM capabilities (Snell et al., 2024; Muennighoff et al., 2025), the increasing time and computational cost on the rollout side are becoming a fundamental performance bottleneck in scaling on-policy RL training (Schulman et al., 2017; Rafailov et al., 2023; Mroueh et al., 2025; Yu et al., 2025b).
@@
Thereby, this raises a critical question: How can we optimize rollouts to reduce inference overhead, especially under long-tail rollout generation in RL training? To address this challenge, we propose Active Partial Rollouts (APRIL), a compute-efficient method to accelerate rollout generation in RL training. In each iteration, we deliberately over-provision rollout requests (exceeding the default batch size) to the inference engines. Once the default number of rollouts is reached, the engines actively stop the remaining unfinished rollouts. Instead of discarding these partial results, we store them in a buffer and resume their generation in the next iteration. Meanwhile, the completed rollouts are used to compute the loss and optimize the LLM. This approach not only safeguards model accuracy by ensuring that no rollout instances are discarded, but also systematically recycles incomplete rollouts to alleviate the long-tail effect, thereby substantially reducing GPU idle time during inference. As a result, throughout the RL training process, APRIL markedly enhances overall training efficiency while preserving the original accuracy. In practice, our experiments demonstrate that APRIL not only delivers at least a 20% improvement (3.2.1) in rollout throughput across widely used RL algorithms—GRPO (Shao et al., 2024), DAPO (Yu et al., 2025b), and GSPO (Zheng et al., 2025)—as well as various LLMs (DeepSeek-AI et al., 2025; Team, 2025), but also achieves faster convergence and higher final accuracy (3.2.2). Moreover, we undertake additional efforts to support deployment across hardware platforms such as NVIDIA and AMD GPUs (3.3).
@@
  • We are the first team to propose and open-source the active partial rollout, APRIL, which improves at least 20% rollout throughput across widely used RL algorithms (3.2.1).
  • In practice, our experiments demonstrate that APRIL not only improves the efficiency of RL, but achieves faster convergence and increases the final accuracy by around 2%5% (3.2.2).
@@ 2.1 Preliminary about General RL Training
$\theta_{k+1}\leftarrow\theta_{k}+\mu\cdot\mathbb{E}_{\underbrace{a\sim{\pi}(\theta_{k})}_{\text{rollout}}}\Big[R(a)\cdot\underbrace{\nabla_{\theta}\log{\pi}(a,\theta_{k})\big|_{\theta=\theta_{k}}}_{\text{training}}\Big],$ (1)
@@ 2.2 Partial Rollout for Mitigating the Long-Tail Problem
In the standard synchronous RL training paradigm, given a batch of instances, the inference engines generate rollouts for all instances ($a\sim\pi(\theta_{k})$) until completion. The resulting trajectories are then forwarded to the training engine ($\nabla_{\theta}\log\pi(a,\theta_{k})|_{\theta=\theta_{k}}$) to update the policy model, as illustrated in formula 1 and Figure 2. However, this procedure often leads to suboptimal GPU utilization, since faster rollouts must wait for the longest sequences to finish, thereby introducing idle time and slowing down the overall training cycle.
@@
To address this inefficiency, we propose the APRIL mechanism as shown in Figure 2. Rather than waiting for all generation workers to complete their sequences, APRIL over-provisions instances ($N^{\prime}$) to the inference engines and terminates generation once the required threshold (batch size $N$) is satisfied at step $k$. The completed rollouts will be used to optimize the policy model ($\pi_{\theta_{k}}$), and uncompleted rollouts are stored and resumed in the subsequent step $k+1$, ensuring no data is wasted while significantly reducing GPU idle time. The process is as follows:
  1. 1.
    Over-provisioned Generation: At the beginning of a rollout phase at step $k$, we initiate $N^{\prime}$ instances where $N^{\prime}>N$.
  2. 2.
    Early Termination: As soon as the rollouts of $N$ instances are fully generated, we dispatch an abort signal to the inference engine, which terminates the generation immediately.
  3. 3.
    Buffering Continuations: These aborted, uncompleted rollout sequences are not discarded. Instead, they are stored in a buffer. The completed sequences are sent to the training engine for optimizing the policy model parameters ($\theta_{k}$).
  4. 4.
    Prioritized Resumption: In the next rollout phase (step $k+1$), the generation workers prioritize resuming these uncompleted rollout sequences from the continuation buffer before starting new ones from scratch.
APRIL approach significantly improves system throughput by eliminating delays caused by the completion of ”long-tail” rollout generations.
@@
  • Newly-completed rollout instances ($\mathcal{D}_{k+1~(completed)}$): Trajectories that were initiated and completed entirely under the current policy $\theta_{k+1}$}.
  • Continued rollout instances ($\mathcal{D}_{k~(completed)}$): Trajectories was initialized by an earlier policy $\pi_{\theta_{k}}$, then aborted and resumed one or more times, and ultimately completed under the current policy $\theta_{k+1}$}.
Consequently, the collected rollout dataset $N$ in $\mathcal{D}_{k+1}$ is not monolithic. Instead, it comprises data generated not only from $\theta_{k+1}$} but also from earlier policies ${\theta_{k}},{\theta_{k-1}},\ldots,{\theta_{k-m}}$, where $m\leq k$.
Therefore, the advantage estimate associated with such trajectories is calculated over a sequence generated by a “mixed” policy. Our empirical results demonstrate that this modification does not destabilize training and occasionally yields modest but consistent gains, as shown in Figure 5. Our method has been successfully applied to multiple variants of PPO, including GRPO (Mroueh et al., 2025), DAPO (Yu et al., 2025a), and GSPO (Zheng et al., 2025).

### v3

@@ Abstract
Reinforcement learning (RL) has become a cornerstone in advancing large-scale pre-trained language models (LLMs). Successive generations—including GPT-o series, DeepSeek-R1, Kimi-K1.5, Grok 4, and GLM-4.5—have relied on large-scale RL training to enhance reasoning and coding capabilities. To meet the community’s growing RL needs, numerous RL frameworks have been proposed. Most of these frameworks primarily rely on inference engines for rollout generation and training engines for policy updates. However, RL training remains computationally expensive, with rollout generation accounting for more than 90% of total runtime. Besides, its efficiency is often constrained by the long-tail distribution of rollout response lengths, where a few lengthy responses stall entire batches, leaving GPUs idle and underutilized. As model and rollout sizes continue to grow, this bottleneck increasingly limits scalability. To address this challenge, we propose Active Partial Rollouts in Reinforcement Learning (APRIL), which mitigates long-tail inefficiency. In the rollout phase, APRIL over-provisions rollout requests, terminates once the target number of responses is reached, and recycles incomplete responses for continuation in future steps. This strategy ensures that no rollouts are discarded while substantially reducing GPU idle time. Experiments show that APRIL improves rollout throughput by 22.5% on average across commonly used RL algorithms (GRPO, DAPO, GSPO), accelerates convergence, and achieves 2.1% higher final accuracy across tasks on average. Moreover, APRIL is both framework- and hardware-agnostic, already integrated into the slime111slime repository: https://github.com/THUDM/slime RL framework, and deployable on NVIDIA and AMD GPUs. Taken together, this work unifies system-level and algorithmic considerations in proposing APRIL, with the aim of advancing RL efficiency and inspiring further optimizations in RL systems. Our codebase is available at: https://github.com/RLsys-Foundation/APRIL
@@ 1 Introduction
This need has been actively addressed by the open-source community, resulting in scalable RL training frameworks such as OpenRLHF (Hu et al., 2025), verl (Sheng et al., 2025), Areal (Fu et al., 2025), and slime (THUDM, 2025). These frameworks typically integrate advanced distributed engines to enhance efficiency—employing vLLM (Kwon et al., 2023) or SGLang (Zheng et al., 2024) as inference backends for LLM rollouts, FSDP (Zhao et al., 2023) or Megatron-LM (Shoeybi et al., 2019) as training backends for LLM optimization, and Ray (Moritz et al., 2018) for orchestrating parallel training and inference to maximize resource utilization during RL training.
Despite extensive infrastructure-level optimizations, the rollout phase remains the dominant bottleneck: each input requires the LLM to generate $\mathcal{N}$ responses of length up to $\mathcal{L}$ in an auto-regressive manner, which typically accounts for over 90% of the total RL training runtime (Hu et al., 2025). Besides, the response length of the rollouts for each input instance often varies significantly in most tasks (3.2). As a result, in any given training batch, a small number of examples may require exceptionally long responses. This ”long-tail” distribution causes the entire batch to stall until the longest-running instance rollout is completed, leading to severe inefficiency and under-utilization of computational resources, as faster-generating workers remain idle (Zhong et al., 2025c). Moreover, with recent findings and the emerging trend of scaling up rollout inference ($\mathcal{N}$ and $\mathcal{L}$) to enhance LLM capabilities (Snell et al., 2024; Muennighoff et al., 2025), the increasing time and computational cost on the rollout side are becoming a fundamental performance bottleneck in scaling on-policy RL training (Schulman et al., 2017; Rafailov et al., 2023; Mroueh et al., 2025; Yu et al., 2025b).
@@
Thereby, this raises a critical question: How can we optimize rollouts to reduce inference overhead, especially under long-tail rollout generation in RL training? To address this challenge, we propose Active Partial Rollouts (APRIL), a compute-efficient method to accelerate rollout generation in RL training. In each iteration, we deliberately over-provision rollout requests (exceeding the default batch size) to the inference engines. Once the default number of rollouts is reached, the engines actively stop the remaining unfinished rollouts. Instead of discarding these partial results, we store them in a buffer and resume their generation in the next iteration. Meanwhile, the completed rollouts are used to compute the loss and optimize the LLM. This approach not only safeguards model accuracy by ensuring that no rollout instances are discarded but also systematically recycles incomplete rollouts to alleviate the long-tail effect, thereby substantially reducing GPU idle time during inference. As a result, throughout the RL training process, APRIL markedly enhances overall training efficiency while preserving the original accuracy. In practice, our experiments demonstrate that APRIL not only delivers a significant improvement (3.2.1) in rollout throughput (up to 49.5% at most) across widely used RL algorithms—GRPO (Shao et al., 2024), DAPO (Yu et al., 2025b), and GSPO (Zheng et al., 2025)—as well as various LLMs (DeepSeek-AI et al., 2025; Team, 2025), but also achieves faster convergence and higher final accuracy (3.2.2). Moreover, we make additional efforts in GPU memory management to support deployment across hardware platforms, such as NVIDIA and AMD (3.3).
@@
  • We are the first to propose and open-source Active Partial Rollout (APRIL), which improves rollout throughput by 22.5% on average across widely algorithms in RL training (3.2.1).
  • In practice, our experiments demonstrate that APRIL not only improves the RL efficiency, but achieves faster convergence and increases the final accuracy by around $\textbf{2.1}\%$ (3.2.2).
@@ 2.1 Preliminary about General RL Training
$\theta_{k+1}\leftarrow\theta_{k}+\mu\cdot\mathbb{E}_{\underbrace{a\sim{\pi}(\theta_{k})}_{\text{Inference: rollout}}}\Big[R(a)\cdot\underbrace{\nabla_{\theta}\log{\pi}(a,\theta_{k})\big|_{\theta=\theta_{k}}}_{\text{Training}}\Big],$ (1)
@@ 2.2 Partial Rollout for Mitigating the Long-Tail Problem
In the standard synchronous RL training paradigm, given a batch of instances, the inference engines generate rollouts for all instances ($a\sim\pi(\theta_{k})$) until completion. The resulting trajectories are then forwarded to the training engine ($\nabla_{\theta}\log\pi(a,\theta_{k})|_{\theta=\theta_{k}}$) to update the policy model, as illustrated in Formula 1 and Figure 2. However, this procedure often leads to suboptimal GPU utilization, since faster rollouts must wait for the longest sequences to finish in the inference engine, thereby introducing idle time and slowing down the overall training cycle.
@@
To mitigate inefficiency, we propose the APRIL mechanism as shown in Figure 2. Instead of waiting for all workers to finish rollouts in the inference stage, APRIL launches an over-provisioned number of instance rollouts ($N^{\prime}>N$) and terminates once the required $N$ instance rollout sequences are completed at step $k$. The finished rollouts are used to optimize the policy model $\pi({\theta_{k}})$, while the unfinished ones are buffered and resumed in the next step $k+1$, reducing GPU idle time without data loss. The process is as follows:
  • Over-provisioned Generation: Start $N^{\prime}$ instances where $N^{\prime}>N$.
  • Early Termination: Stop generation once $N$ instance rollout sequences are completed.
  • Buffering Continuations: Store unfinished instance rollout rollouts in a continuation buffer; completed ones are sent for policy optimization.
  • Prioritized Resumption: At step $k+1$, resume buffered rollouts before starting new ones.
APRIL approach significantly improves system throughput by eliminating delays caused by the completion of ”long-tail” rollout generations. However, an important point should be noted next.
@@
  • Newly-completed rollout instances ($\mathcal{D}_{k+1~(completed)}$): Trajectories that were initiated and completed entirely under the current policy $\pi({\theta_{k+1}})$.
  • Continued rollout instances ($\mathcal{D}_{k~(completed)}$): Trajectories was initialized by an earlier policy $\pi({\theta_{k}})$, then aborted and resumed one or more times, and ultimately completed under the current policy $\pi({\theta_{k+1}})$.
Consequently, the collected rollout dataset $N$ in $\mathcal{D}_{k+1}$ is not monolithic. Instead, it comprises data generated not only from $\pi({\theta_{k+1}})$ but also from preceding policies $\pi({\theta_{k}}),\pi({\theta_{k-1}}),\ldots,\pi({\theta_{k-m}})$, where $m\leq k$.
Therefore, the advantage estimate associated with such trajectories is calculated over a sequence generated by a “mixed” policy. Our empirical results demonstrate that this modification does not destabilize training and occasionally yields modest but consistent gains, as shown in Figure 7. Our method has been successfully applied to multiple most common used RL algorithms, including PPO (Schulman et al., 2017), GRPO (Mroueh et al., 2025), DAPO (Yu et al., 2025a), and GSPO (Zheng et al., 2025).